Thanks for your help. I am a mom trying to help my daughter and we both are stuck on this one. Please see if you can help us understand it. The problem is from a worksheet. The question is: Find the measure of an angle for which the sum of the measures of its complement and its supplement are 178 degrees.
Answer by ankor@dixie-net.com(12688)

he question is: Find the measure of an angle for which the sum of the measures of its complement and its supplement are 178 degrees.
:
Let the angle = x
:
Complement 90 - x
Supplement: 180 - x
:
Complement + supplement = 178
(90 - x) + (180 - x) = 178
90 + 180 - x - x = 178
270 - 2x = 178
-2x = 178 - 270
-2x = -92
x = -92/-2
x = + 46 degrees
:
Did this help?